### v1

11institutetext: Department of Physics and Astronomy, University of California, Los Angeles, CA 90095, USA; 11email: pmozumdar@astro.ucla.edu 22institutetext: Department of Astronomy, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China 33institutetext: Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai Jiao Tong University, Shanghai 200240, China 44institutetext: Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education, Shanghai Jiao Tong University, Shanghai 200240, China 55institutetext: Department of Astronomy & Astrophysics, University of Chicago, Chicago, IL 60637, USA 66institutetext: Kavli Institute for Cosmological Physics, University of Chicago, Chicago, IL 60637, USA 77institutetext: Center for Astronomy, Space Science and Astrophysics, Independent University, Bangladesh, Dhaka 1229, Bangladesh 88institutetext: Sub-Department of Astrophysics, Department of Physics, University of Oxford, Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK 99institutetext: Dipartimento di Fisica e Astronomia “Augusto Righi”, Alma Mater Studiorum - Università di Bologna, via Gobetti 93/2, 40129, Bologna, Italy
TDCOSMO
XXII. Accurate stellar velocity dispersions of the SL2S lens sample and the lensing mass fundamental plane.
[…]
(Received xxx, xxxx; accepted xxx, xxxx)
We reanalyze spectra taken as part of the SL2S lens galaxy survey, with the goal to obtain stellar velocity dispersion with precision and accuracy sufficient for time-delay cosmography. In order to achieve this goal, we impose stringent cuts on signal-to-noise ratio (SNR), and employ recently developed methods to mitigate and quantify residual systematic errors due to template libraries and fitting process. We also quantify the covariance across the sample. For galaxies with spectra with SNR >20/>20/> 20 /Å, our new measurements have average random uncertainty of 3-4%, average systematic uncertainty of 2%, and covariance across the sample of 1%. We find negligible covariance between spectra taken with different instruments. The systematic uncertainty and covariance need to be included when the sample is used as an external dataset in time-delay cosmography. We revisit empirical scaling relations of lens galaxies based on the improved kinematics. We show that the SLS2 sample, the TDCOSMO time-delay lens sample, and the lower redshift SLACS sample follow the same correlation between effective radius, stellar velocity dispersion and lensing mass, known as the lensing mass fundamental plane, as that derived by Auger et al. (2010) assuming isothermal mass profiles for the deflectors. We also derive for the first time the lensing mass fundamental plane assuming free power-law mass density profiles, and show that the three samples also follow the same correlation. This is consistent with a scenario in which massive galaxies evolve by growing their radii and mass, but staying within the plane.
Key Words.:
cosmology: distance scale – gravitational lensing: strong – Galaxy: kinematics and dynamics – Galaxies: elliptical and lenticular, cD
1 Introduction
The Hubble tension is arguably one of the most pressing open problems in cosmology (e.g., Di Valentino et al. 2025, and references therein.). The tension between the Hubble Constant H0 inferred from early and late universe probes has surpassed the canonical 5-$\sigma$ threshold for a statistically significant discrepancy. If the tension is not due to yet unidentified systematic errors, it implies that new physics is required, beyond the canonical flat $\Lambda$CDM.
In order to quantify whether the tension is due to systematic errors, methods to measure H0 independent of the traditional local distance ladder and the sound horizon at recombination are especially valuable. Among these, the so-called time-delay cosmography stands out (Refsdal 1964; Treu et al. 2016, 2022; Birrer et al. 2024). It is completely independent of all other methods, it measures distances that are large enough to be unaffected by local over/underdensities and peculiar motions, and yet are small enough to be regarded as a late-time probe.
State-of-the-art time-delay cosmography relies on stellar velocity dispersion measurements of the galaxies acting as deflectors (Treu & Koopmans 2002; Koopmans et al. 2003; Shajib et al. 2018; Yıldırım et al. 2020; Birrer et al. 2020; Shajib et al. 2023) to break the mass-sheet degeneracy (Falco et al. 1985). However, the requirements for stellar velocity dispersion measurements for time-delay cosmography are much more stringent than for galaxy evolution studies. The relative error on H0 is approximately double the relative error on stellar velocity dispersion, and thus one needs to control systematic errors and covariances between measurements at the percent level or better in order to achieve the 2% precision and accuracy required to distinguish between the Planck (Planck Collaboration et al. 2020) and SH0ES measurements (Riess et al. 2022). Motivated by this requirement, our team carried out a detailed investigation of stellar velocity dispersion kinematics. Knabel et al. (2025) showed that using carefully curated stellar template libraries and methods it is possible to achieve sub-percent accuracy, if spectra of sufficient SNR and resolution are available.
In this paper, we re-analyze the spectra taken a decade ago as part of the SL2S survey (Ruff et al. 2011; Gavazzi et al. 2012; Sonnenfeld et al. 2013a, b, 2014, 2015) with the goal of implementing the recent improvements in the derivation of stellar kinematics, obtain accurate measurements and estimates of systematic errors and covariance between galaxies. SL2S is a very valuable external dataset for cosmography as it is well matched to the time-delay lenses in terms of redshift distribution (Birrer & Treu 2021). However, while 5-10% precision obtained by Sonnenfeld et al. (2013a) was sufficient to meet the main goal of the SL2S papers, i.e. to characterize the cosmic evolution of luminous and dark matter in early-type galaxies (Sheu et al. 2025, submitted), using them as external datasets for cosmography requires additional effort and more modern techniques.
We also take advantage of the new measurements to investigate whether the SL2S sample galaxies obey the same lensing mass fundamental plane as the lower-redshift SLACS sample (Bolton et al. 2007, 2008b; Auger et al. 2010). The lensing mass plane is an empirical correlation similar to the traditional fundamental plane (Dressler et al. 1987; Djorgovski & Davis 1987), replacing surface brightness with total surface mass density obtained via the lens model. The lensing mass fundamental plane is tighter than the traditional fundamental plane and closer to the simple ”virial” relation, indicating a high degree of similarity between the mass distribution and kinematic structure of massive elliptical galaxies (Nipoti et al. 2008). We use the canonical lensing fundamental plane based on the assumption of isothermal mass density profile, and we introduce for the first time a new plane based on the lensing-only slope measurements to extrapolate the surface mass density to the standard circular aperture of radius equal to half the effective radius. We show that the SL2S and SLACS lenses follow the same planes as the time-delay lenses of the TDCOSMO sample (Millon et al. 2020; Birrer et al. 2020, and TDCOSMO collaboration 2025, in preparation). This finding is consistent with the assumption that external datasets have the same mass-structure as time-delay lenses necessary for obtaining joint constraints (Birrer et al. 2020; Birrer & Treu 2021).
In Section 2, we summarize the SL2S spectroscopic data. In Section 3 we describe our methodology. In Section 4 we present our new measurements of stellar velocity dispersions and their uncertainties. In Section 5 we compare our new measurements with the previously published ones, and we study the lensing mass fundamental plane of the SL2S sample in comparison with SLACS and TDCOSMO. In Section 6, we summarize and discuss our results. When necessary, we adopt a flat $\Lambda$ cold dark matter cosmology, with $\Omega_{\rm m}=0.3$, $\Omega_{\Lambda}=0.7$ and $h=0.7$.
2 SL2S spectroscopic data
Spectroscopy for 58 SL2S candidates was obtained over several years with the goals of confirming the lensing hypothesis, securing source and deflector redshift, and measuring stellar velocity dispersion. Among these candidates, five were confirmed as non-lens, and the other 53 were analyzed by Sonnenfeld et al. (2013b). In this work, we excluded 6 of these 53 systems due to insufficient instrumental resolution and/or very low SNR, as described below in more detail. If multiple spectra were available for a galaxy from the same or different instrument, then we used the one with the highest SNR as our fiducial spectrum. We report also the results from the second-best spectra for completeness and we use them to perform a test of systematic errors. In this section, we briefly describe the three telescope/instrument setups that yielded the spectra of the galaxies in this sample. For specific details of the observations such as observing date, slit width, total exposure, and seeing, etc., check Table 2 in Sonnenfeld et al. (2013b).
2.1 Keck-LRIS
Among the 47 lens galaxies in this sample, 31 were observed with the Low-Resolution Imaging Spectrometer (LRIS; Oke et al. 1995) on Keck. LRIS consists of two detectors (blue and red) covering a wavelength range of 3000-10000 Å. The spectral resolution and reciprocal dispersion depend on the configuration of the grism (blue side) and grating (red side) used. The spectra were collected in the long-slit mode using both arms simultaneously with different exposure times. The observing nights were spread over several years, from 2006 to 2012, with typical seeing conditions (FWHM) ranging $\sim$ 0.6 ′′ - 1.3 ′′. In this paper, we only used spectra from the red side as the instrumental resolution for the setup used on the blue side was very low, prioritizing wavelength coverage for redshift determination. The 31 spectra with sufficient spectral resolution and SNR to determine stellar velocity dispersions are shown in Figures 1 and 2.
Figure 1: Spectra of the 16 galaxies from the 31 observed with LRIS on the Keck Telescope. In each panel, the black line is the observed spectrum, and the red line is the best-fit model from pPXF. The gray regions mark the excluded wavelength range from fitting and typically correspond to atmospheric absorption lines. Each panel also indicates the name of the system, redshift of the galaxy, and prominent absorption features present in the spectra.
Figure 2: As figure 1 for the remaining 15 of the 31 galaxies observed with LRIS.
2.2 VLT-Xshooter
The Xshooter spectrograph on VLT (Vernet et al. 2011) was used to obtain data for 13 systems. Xshooter consists of three independent cross-dispersed echelle spectrographs in three bands (UVB, VIS, NIR) and covers a huge wavelength range of 3000 - 25000 Å. This instrument has a higher resolution than LRIS and DEIMOS. The observations were performed using long slits of width 1.0′′ for the UVB arm, and 0.9′′ for the VIS and NIR arm. The observing nights were spread over three years from 2010 - 2012. The 13 spectra with sufficient spectral resolution and SNR to determine stellar velocity dispersions are shown in Figure 3.
Figure 3: Spectra of 13 galaxies observed with the Xshooter spectrograph on the VLT. In each panel, the gray line shows the observed spectrum, and the red line presents the best-fit model from pPXF. We have also plotted the smoothed version (the yellow line) of the observed spectrum. The gray bands mark the excluded wavelength range from fitting and typically correspond to the atmospheric absorption lines. Each panel also indicates the name of the system, redshift of the galaxy, and prominent absorption features present in the spectra.
2.3 Keck-DEIMOS
Three systems in this sample have data obtained with DEIMOS at Keck Observatory (Faber et al. 2003). The observations were performed using a long slit of width 1′′ with 600ZD grating. This setup covers a wavelength range of 4500 - 9500 Å with reciprocal dispersion of 0.65 Å. The typical instrumental resolution is around $\sigma_{\rm inst}\sim$ 70 km s-1.The spectra were collected during two observing nights in 2012. The 3 spectra with sufficient spectral resolution and SNR ratio to determine stellar velocity dispersions are shown in Figure 4.
Figure 4: Spectra of 3 galaxies from the SL2S sample observed with DEIMOS on the Keck Telescope. In each panel, the black line is the observed spectrum, and the red line is the best-fit model from pPXF. The gray regions mark the excluded wavelength range from fitting.
3 Methodology for new kinematic measurements
3.1 Summary of the method
We apply the recipe proposed by Knabel et al. (2025, hereafter KM25) to measure stellar kinematics and minimize and estimate the associated systematic uncertainties. First, we measured the instrumental line spread function (LSF), $\sigma_{\text{inst}}$ of the observed galaxy spectra. We also measured the SNR of each spectrum and, using simulations, determined the SNR required for each specific instrument to keep any bias below the 1% level. Using the information from KM25, we constructed two clean stellar libraries – Indo-US [^1] https://noirlab.edu/science/observing-noirlab/observing-kitt-peak/telescope-and-instrument-documentation/cflib (Valdes et al. 2004), and MILES [^2] https://miles.iac.es/ (Sánchez-Blázquez et al. 2006; Falcón-Barroso et al. 2011), customized for the wavelength range used in the kinematics fitting, as described in more detail below.
Second, improper flux calibration, variance in continuum strength, etc., may bias the fit significantly if not corrected for using additive and multiplicative polynomials. To find optimal orders of the polynomials and strike a balance between under- and overfitting, we ran a grid of additive and multiplicative polynomials. We find that for low polynomial orders the fits are generally unstable, leave significant residuals, and result in disagreements between the template libraries, likely arising from improper flux calibration. Above a certain threshold, which depends on the wavelength range and on the quality of the flux calibration, the result becomes stable and yields good agreement between the templates. If the polynomial order is too high, the polynomials start modifying the shape of the absorption lines, and the fit becomes unstable again. So, it is important to choose polynomial orders in the ”sweet spot”, as discussed by Knabel et al. (2025). For this dataset, we find that orders of 8 and 2 for additive and multiplicative polynomials respectively, provide the least systematic scatter between templates, although similar results are found for somewhat lower and higher orders.
Third, we select the wavelength range to fit. We intend to use the maximum possible wavelength range available for a galaxy spectrum to maximize the information content without compromising the SNR.
Fourth, following the KM25 recipe, we use the Bayesian information criterion (BIC) to estimate the weighted average velocity dispersion ($\bar{\sigma}$) and associated systematic ($\Delta\bar{\sigma}$) and statistical uncertainties ($\delta\bar{\sigma}$). An important step in accomplishing this goal is to measure the threshold level for stellar templates to be considered as a significant contribution to the fit. If the threshold level is 1%, that means the templates with weight (w) $0.010.0<w<$ 0.01 are not contributing significantly to the fit (improving $\chi^{2}$) and should not be considered when counting the number of templates used to create the model. We measured the BIC of the fits for the whole sample, the BIC weight of each stellar library, and the BIC-weighted velocity dispersion and uncertainties using the equations provided by KM25.
Figure 5: Example of typical variation in $\sigma_{\text{inst}}$ as a function of wavelength for three instruments – LRIS-red, Xshooter, and DEIMOS. In the plot for each instrument, there are two panels. The top one showing $\sigma_{\text{inst}}$ in km s-1 unit and the bottom one in Å unit over observed wavelength. In all panels the black line shows the average $\sigma_{\text{inst}}$ over the wavelength range.
3.2 Determination of instrumental resolution and threshold
We measured the instrumental resolution ($\sigma_{\text{inst}}$) or LSF for each galaxy using its respective variance spectrum. In this regard, we fitted Gaussian profiles plus a constant (for background noise) to the isolated and unblended sky lines at different wavelengths. The fitted second moment of a Gaussian profile is considered $\sigma_{\text{inst}}$ at that wavelength. The $\sigma_{\text{inst}}$ is first measured in Å and then converted to km s-1 using $\sigma_{\text{inst}}=\frac{\Delta\lambda c}{\lambda}$. In Figure 5, we show how $\sigma_{\text{inst}}$ varies typically as a function of wavelength for three instruments – LRIS-red, Xshooter, and DEIMOS. As $\sigma_{\text{inst}}$ for all three instruments does not change significantly over the fitted wavelength range, we use the respective average $\sigma_{\text{inst}}$ within the fitted wavelength range of each galaxy for kinematics measurements. We report the average $\sigma_{\text{inst}}$ of each galaxy in Table LABEL:table:veldis_measurements. The associated uncertainties on $\sigma_{\text{inst}}$ are the sample standard deviations of the measured $\sigma_{\text{inst}}$ at different wavelengths.
We impose a criterion to ensure that any bias due to imprecise measurement of $\sigma_{\text{inst}}$ is less than 1%. Following KM25 this requires:
$\frac{\delta\sigma}{\sigma}\approx\frac{\delta\sigma_{\text{inst}}}{\sigma_{\text{inst}}}\left(\frac{\sigma_{\text{inst}}}{\sigma}\right)^{2}\lesssim 0.01.$ (1)
Thus, depending on the measured $\sigma$, we can only accept instrumental setups for which the criterion is satisfied. We found that all 47 galaxies listed in Table LABEL:table:veldis_measurements satisfy this requirement.
3.3 Determination of SNR threshold
It is well known that spectra with insufficient SNR could result in biased kinematics (e.g., Treu et al. 2001). To determine the SNR required for the unbiased measurements below 1%, we performed numerical simulations with each specific instrumental setup (resolution, sampling, etc.) used to collect the galaxy spectra of the SL2S sample. We conducted simulations for five setups: LRIS-blue, LRIS-red low-resolution, LRIS-red high-resolution, DEIMOS, and Xshooter. We started by creating mock galaxy spectra using G/K-type stars from the Indo-US and MILES stellar library. We resampled the selected template spectra linearly as per the reciprocal dispersion of a particular instrument setup and convolved the spectra with a Gaussian kernel using pPXF. The $\sigma$ value of the Gaussian kernel is determined by the velocity dispersion to be injected and the instrumental resolution, $\sigma_{\text{inst}}$, of the setup. Then we added noise to the convolved spectra to mimic a particular SNR. The reciprocal dispersions we used for these setups are 1.45, 1.85, 0.8, 0.65, 0.2 Å respectively. Similarly, the instrumental resolutions applied on the mock galaxy spectra for the setups are $\sigma_{\text{inst}}=$ 200, 200, 80, 70, 35 km s-1 respectively. The chosen reciprocal dispersions and $\sigma_{\text{inst}}$ are representative of the actual SL2S data collected by each setup. For a range of velocity dispersions values ( $\sigma_{\text{input}}$ = 150, 200, 250, 300 km s-1) and SNR (10, 15, 20, 25, 30, 35, 40), we created 1000 mock galaxy spectra and recovered the input velocity dispersion using clean Indo-Us stellar library and the pPXF package [^3] https://pypi.org/project/ppxf/ (Cappellari 2017, 2023).
From these simulations, we found that, for LRIS-red high-resolution, DEIMOS, and Xshooter, a SNR threshold of 15 per Å is sufficient to obtain an unbiased estimate below 1%. However, LRIS-blue and LRIS-red low-resolution setups do not have sufficient spectral resolution to guarantee bias below 1% even at high SNR and velocity dispersion. Additionally, as the $\sigma_{\text{inst}}$ of these two setups are comparable to the velocity dispersions of the galaxies to be measured ($\sim$ 200 km s-1), we decided to exclude the spectra from these two setups in this paper. Fortunately, we have to discard only two systems from the sample due to this constraint. In all other cases there are complementary data available from the LRIS-red high-resolution setup.
3.4 Measuring SNR of the spectra
We measured the SNR per Å of all available spectra from the three accepted instrument setups described in the last section. We used three different rest frame wavelength intervals — 4000–4400, 4800–5200, and 5400–5800 Å, to determine the average SNR per pixel of the spectra and divided it by the square root of the number of Å per pixel to obtain the average SNR per Å. We selected these three wavelength regions as they have relatively uniform SNR per pixel. If a spectrum does not contain all three regions, then we use whatever wavelength ranges from these three regions are available. We excluded pixels covering atmospheric A and B lines if present in these three regions while measuring average SNR. We found 30 galaxies with spectra having SNR per Å $\geq$ 15. However, we still measured and reported kinematics and associated uncertainties of the other galaxies as they could be used in projects where bias above 1% is acceptable. The SNR of the fitted galaxy spectra is listed in Table LABEL:table:veldis_measurements.
3.5 Selection of clean stellar libraries
KM25 discussed extensively the importance of using clean stellar templates and showed that the usage of defective templates can introduce bias in the kinematics measurements. Therefore, we have created two clean stellar libraries– Indo-Us and MILES using the list of clean stars provided in KM25. Some of the stars in these lists also contain defects in several wavelength regions, and KM25 further flagged these stars with the information of the problematic wavelength range. We checked for the presence of these defective regions in the clean templates within the wavelength range of 3600-6000 Å. If defects are present, we also excluded these spectra from the clean library. We used the clean Indo-US and MILES library for all the kinematics measurements and systematic tests. The final clean Indo-US and MILES libraries contain 802 and 789 templates, respectively.
We initially also considered using the Xshooter stellar library. However, more than 50% of the clean star templates in this library have been flagged beyond 5300 Å due to artifacts around the dichroic separating the UVB from the visible wavelength range, while for most galaxies we have to fit up to 6000 Å at rest-frame. As the clean Xshooter library has to be downsized significantly beyond 5300 Å, we decided not to use this library in this project.
3.6 Determination of polynomial orders
Following the recipe in KM25, we ran a grid of polynomial orders to determine the optimal orders of the additive and multiplicative polynomials that would make the kinematics measurements stable and reduce the systematic scatter in measurements using different template libraries for the whole sample. A similar test was conducted by KM25 using a sample of galaxy spectra covering the same rest-frame wavelength range and observed with the same instrument. However, the galaxy spectra in the SL2S sample contain different rest-frame wavelength ranges and were collected using different instruments. Therefore, we also checked whether the optimal polynomial orders depend on the instrument or the fitted wavelength range. We found that higher polynomial orders, such as 7-9 for additive polynomials and 2 for multiplicative polynomials, reduce the systematic scatter regardless of the instrument or wavelength range. We decided to use orders of 8 and 2, for additive and multiplicative polynomials, respectively. As a test, we also computed what would be the additional scatter in stellar velocity dispersion if we marginalized over polynomial orders around the minimal choice, and we found it to be much less than 0.5%, therefore negligible with respect to that attributable to stellar templates.
3.7 Wavelength range and masking
The galaxy spectra of the SL2S sample cover different rest-frame wavelength ranges due to combination of redshift and instrument setup. Regardless, we used the maximum possible wavelength range available for a galaxy spectrum without compromising the SNR. The cutoff of the adopted wavelength range on the blue side is imposed both by the instrument setup and stellar template coverage, and on the red side by the redshifts of the galaxies. In Figures 1 - 4, we show the fitted wavelength range for each galaxy. We excluded some wavelength regions from the fits at two stages. First, we masked the rest-frame wavelength range 6850-6950 Å and 7580-7750 Å  corresponding to the atmospheric B-band and A-band, respectively, if present. Next, using an iterative sigma clipping method, we detected wavelength ranges containing emission lines, instrument or extraction related artifacts, cosmic rays, etc., if any, and also masked them from the fit. In Figures 1 - 4, these masked regions are shown in gray bands.
[…]
As an additional test, we repeated the measurement for a subset of galaxies with spectra with SNR $>20$/Å, masking at random one of the most prominent absorption features present in the spectra (CaHK, Mgb, G-band, NaD). The results are shown in Figure 6, separately for Indo-US and MILES library, in order to isolate the effect of the lines from that of the stellar library. On average the stellar velocity dispersion changes by less than 1%, demonstrating that for data and templates of sufficient quality the results do not depend on any single feature.
Figure 6: For each galaxy with high-quality spectra in the sample (SNR$>20$/Å) we show the mean and scatter in the value of stellar velocity dispersion when masking one of the main absorption features (CaHK, G-band, MgB, NaD) with respect to the value obtained for the entire spectrum. Thick horizontal lines represent the sample average and dashed lines the error on the mean.
3.8 Determining the threshold to count effective template number
For the reasons discussed earlier, we used two clean stellar libraries for kinematic fitting and combined the measurements and associated uncertainties as a weighted average. The weight assigned to each template library is determined using BIC for the whole sample. The total BIC and associated weight is calculated following equations 5-10 in KM25. One crucial step in measuring BIC is to determine the effective number of templates that provide a significant contribution to the kinematic model. KM25 extensively discussed the issue and how to address it. Briefly, in an example of overfitting the noise, the model could include some templates with very low weights which may improve the fit (lowering $\chi^{2}$) by a minimal amount. However, they degrade the BIC significantly by increasing the total number of parameters. This effect can lead to incorrect estimates of the weights for the template libraries. Following the procedure in KM25, we first carried out a pPXF fit with all the templates in a clean library, and then performed successive fits while only using the subset of templates that contribute only above a certain threshold to the initial fit. The threshold is the weight of the templates in defining the model for the initial fit. As expected, we find that for a higher threshold (10-20%), the BIC is relatively large because there are not enough templates to obtain a good fit. However, below $\sim 5$%, the $\chi^{2}$ does not improve enough to compensate for the increased number of degrees of freedom in the BIC. Therefore, we adopted a 5% threshold to count the effective number of templates for the SL2S sample (as expected the threshold is similar to the inverse of the SNR of the dataset, see KM25).
Figure 7: BIC-weighted velocity dispersions of the SL2S sample as a function of redshift. The green, red, and black markers show galaxies in three different SNR bins - SNR ¿ 20, 13 ¡ SNR $\leq$ 20, and SNR ¿20, respectively. The associated error bars show total uncertainties in the measured velocity dispersions and are dominated by the statistical uncertainties, which are driven by the SNRs.
4 New measurements and uncertainties
In this section, we present the results obtained by using the methodology discussed in the previous section. The fitted models from pPXF are shown in Figures 1 to 4, and new measurements and uncertainties are summarized in Table LABEL:table:veldis_measurements. For each galaxy, we report the SNR and instrumental resolution of the spectrum, our best estimate of the stellar velocity dispersion, together with the random and systematic errors. The mean $\bar{\sigma}$, the systematic uncertainty $\Delta\bar{\sigma}$, and statistical uncertainty $\delta\bar{\sigma}$ of the velocity dispersion of an individual galaxy are defined as in KM25 such that
$\bar{\sigma}=\frac{\sum_{k}w_{k}\sigma_{k}}{\sum_{k}w_{k}}$ (2)
$(\Delta\bar{\sigma})^{2}=\frac{\sum_{k}w_{k}(\sigma_{k}-\bar{\sigma})^{2}}{\sum_{k}w_{k}},$ (3)
and
$\delta\bar{\sigma}=\sqrt{\frac{\sum_{k}w_{k}(\delta\sigma_{k})^{2}}{\sum_{k}w_{k}}},$ (4)
where the index $k$ runs through the template library, $\delta\sigma_{k}$ is the statistical uncertainty estimated by pPXF, and $w_{k}$ is the weight assigned.
Following KM25, we provide the stellar velocity dispersions and uncertainties as both the BIC-weighted average and equally-weighted average of the measurements from two template libraries. The BIC determined weights were estimated at the sample level to minimize noise in the process, and the samples were grouped by instrument and/or SNR. The estimated weights for the groups are summarized in Table 2. Note that preference for one of the libraries changes depending on the instrument and the SNR. For each galaxy, we used the weight for the group it is affiliated with to measure BIC-weighted average values. For example, if a galaxy is observed with LRIS and has an SNR$>$20/Å, then the assigned weights to the measurements from Indo-US and MILES are 0.78 and 0.22, respectively. In general, the BIC-weighted and equally-weighted velocity dispersions are very similar, consistent with the good agreement between the template libraries.
When the number of template libraries is small (as in the case considered here), it is prudent to adopt Bessel’s correction to get the unbiased estimators (see KM25 for a derivation). Thus, the systematic uncertainties are Bessel corrected following:
$(\Delta_{B}\bar{\sigma})^{2}=\frac{\sum_{k}w_{k}(\sigma_{k}-\bar{\sigma})^{2}}{\sum_{k}w_{k}-\sum_{k}w_{k}^{2}/\sum_{k}w_{k}}.$ (5)
However, for $k=2$, the Bessel correction nullifies the dependence on the weights and $\Delta_{B}\bar{\sigma}$ reduces to $(\Delta_{B}\bar{\sigma})^{2}=(\frac{\sigma_{1}-\sigma_{2}}{2})^{2}$ where $\sigma_{1}$, and $\sigma_{2}$ are measurements using two template libraries. Thus, the Bessel-corrected systematic uncertainties, $\Delta_{B}\bar{\sigma}$, for the BIC-weighted and equally-weighted velocity dispersions are equal and reported as a single column in Table LABEL:table:veldis_measurements.
We have also measured the BIC weighted average statistical and systematic uncertainty, both Bessel corrected and non-corrected, for subsamples grouped by SNR and/or instrumental setup. In addition, we estimate the covariance matrix as follows:
$C_{ij}\bar{\sigma}=\frac{\sum_{k}w_{k}(\sigma_{k,i}-\bar{\sigma}_{i})(\sigma_{k,j}-\bar{\sigma}_{j})}{\sum_{k}w_{k}}+\delta_{i,j}\delta\bar{\sigma}_{i}^{2},$ (6)
where the indices $i,j$ identify the elements of the covariance matrix, $\bar{\sigma}_{i,j}$ is the weighted average velocity dispersion of the element at index $i$ or $j$, and $\delta_{i,j}$ is the Kronecker delta function. And as above, the index $k$ runs through the template library, $\delta\bar{\sigma}_{i}$ is the weighted average statistical uncertainty, and $w_{k}$ is the weight assigned. Along the diagonal, the first term is the systematic error, and the second is the statistical error. With Bessel correction, the estimator reads as
$C_{B,ij}\bar{\sigma}=\frac{\sum_{k}w_{k}(\sigma_{k,i}-\bar{\sigma}_{i})(\sigma_{k,j}-\bar{\sigma}_{j})}{\sum_{k}w_{k}-\sum_{k}w_{k}^{2}/\sum_{k}w_{k}}+\delta_{i,j}\delta\bar{\sigma}_{i}^{2}.$ (7)
The normalized sample average uncertainties and the average amplitude of the off-diagonal terms of the covariance matrix between elements are summarized in Table 2. As expected, the random error is driven by the SNR, regardless of the instrumental setup, and ranges between 3-8%. The systematic errors also depend on the SNR and instrumental setup, and are at the level of 1-2% without Bessel correction and around 2-3% with the correction. This is, of course, only the error associated with template choice. Other potential sources of error, such as the wavelength range and polynomial order, were ignored, as they are negligible with respect to the template contribution.
The average off-diagonal amplitude of the covariance matrix for different subsamples is positive. This is expected as the templates yielding higher (lower) velocity dispersion do so consistently for all galaxies. This quantity ranges around 1-2% within the same instrumental setup, and subpercent across different instrumental setups. The LRIS subsamples have slightly higher covariances than the combined Xshooter and DEIMOS samples.
We found that Indo-US library is preferred by the LRIS data, while MILES is preferred by the other. This emphasizes the importance of using several stellar libraries and calculating BIC-weighted averages. However, in this case, similar results are obtained if the templates are given equal weight, so the difference in BIC weight has less impact.
As a consistency test of our error estimates, we compare the stellar velocity dispersion obtained for 11 systems that have multiple spectra with sufficient resolution. The fiducial spectrum is the one with the highest SNR. The other spectra may have significantly lower SNR and may cover wavelength ranges differing by 500-1000 Å from their fiducial counterpart. The difference in the SNR quality of the spectra is reflected in the random uncertainties listed in Table LABEL:table:veldis_measurements. The comparison is shown in Figure 8, with the error bars representing the sum in quadrature of the random and systematic errors. The sample average of the ratio is $<\frac{\sigma_{\mathrm{high\ SNR}}}{\sigma_{\mathrm{low\ SNR}}}-1>=-0.11\pm 2.70$. Repeated measurements with higher SNRs are needed to carry out an even more stringent test of systematics across different instrumental setups.
[…]
5 Comparison with previous work and improved scaling relations
In this section, we show that the new measurements are consistent with the previous ones (Figure 9), with no average offset. We compared the BIC-weighted velocity dispersion of 37 galaxies with the corresponding measurements in Sonnenfeld et al. (2013b) for which the same spectra were used. For other cases, different spectra were used, or no measurements were reported. We found that for the 37 galaxies, the estimated $<y_{i}/x_{i}>=0.98\pm 0.02$ where $y_{i}$ and $x_{i}$ are new and old measurements for the $i$th galaxy. Thus, the current velocity dispersions are consistent with the previously measured values on average. The reduction in systematic error with respect with the previous analysis is due mostly to three factors. First, stellar template libraries have been improved. Whereas Sonnenfeld et al. (2013b) used 7 stars from the Indo-US library, we use hundreds of stars from the cleaned Indo-US and MILES libraries. Second, we do not use the data taken with instrumental setups that have insufficient spectral resolution to meet our target accuracy. Third, we have mitigated the effects of continuum and flux calibration uncertainties by carefully modeling them with multiplicative and additive polynomials.
The distribution of velocity dispersions in the SL2S sample does not vary significantly with redshift (as we show in Figure 7), and peaks at significantly higher redshift than SLACS ($z=0.1-0.4$).
Figure 9: Comparison of velocity dispersions for 37 galaxies as measured in this paper (y-axis) and by Sonnenfeld et al. (2013a, x-axis) based on the same spectra. The only thing that changed is the methodology, chiefly improvements in the stellar template libraries.
We then use the new SL2S measurements to revisit the scaling relations determined by Auger et al. (2010) for the SLACS sample, taking advantage of the new spectroscopy determinations by Knabel et al. (2024) and those presented here. Note that to enable direct comparison to Auger et al. (2010), the stellar velocity dispersion has been corrected to a fiducial aperture of half the effective radius using the empirical recipes derived by Knabel et al. (2024). We refer to these aperture corrected velocity dispersions as $\sigma_{1/2}$. We adopt total uncertainties as the sum in quadrature of the random and systematic errors.
Our goal is to investigate whether there are any differences between the SLACS, SL2S, and TDCOSMO samples (Millon et al. 2020; Birrer et al. 2020, and TDOCOSMO collaboration 2025, in preparation) that may potentially be an indication of structural evolution or differences in terms of these properties.
We limit ourselves to the SLACS and SL2S galaxies with high SNR spectroscopy from Knabel et al. (2024, 2025) and this work, to minimize the potential impact of systematic errors. We also limit ourselves to the scaling relations that involve lensing quantities, as the others can be investigated using much larger samples of non-lens galaxies. Effective radii and Einstein radii are taken from Auger et al. (2010) for the SLACS sample, from Sheu et al. (2025; submitted) for the SL2S sample, and from TDCOSMO collaboration (2025, in prep) for the TDCOSMO sample. The effective radii are obtained by fitting Sérsic (1963) profiles to the surface brightness of the deflector, jointly with the lensed images444de Vaucouleurs (1948) in the case of SLACS-X. Noting that the formal errors are usually tiny on these quantities, we adopt 10% as our systematic error on the effective radius (e.g., Cappellari et al. 2013a, sec. 4.1.1) and 2% on the Einstein radius (e.g., Bolton et al. 2008a). We neglect possible contributions to the lensing quantities by external convergence, since it is at a few percent level at most (Birrer et al. 2020) and would not affect our conclusions in a significant way.
In addition to the scaling relations described by Auger et al. (2010), we consider a new correlation that involve the mass density profile slope $\gamma_{\rm lens}$ determined from lens models (Tan et al. 2024, and Sheu et al. 2025, submitted). This quantity was not available to Auger et al. (2010). We note that Auger et al. (2010) used a slope $\gamma$ derived by a joint lensing and dynamical analysis, which is the same only if the mass density profiles are exactly power laws (Sonnenfeld et al. 2013b).
All the quantities presented in this analysis are summarized in Table 3.
5.1 Correlation between lensing mass profile slope and ratio between lensing and stellar velocity dispersion
The first correlation that we examine is one between lensing slope and the ratio between the stellar velocity dispersion $\sigma$ and lensing velocity dispersion $\sigma_{\rm SIE}$, $f_{\rm SIE}=\sigma/\sigma_{\rm SIE}$. The lensing velocity dispersion is related to the Einstein radius $\theta_{\rm E}$ and the angular diameter distance to the source (Ds) and between the deflector and the source (Dds) by the equation (Treu 2010):
$\theta_{E}=4\pi\left(\frac{\sigma_{\rm SIE}}{c}\right)^{2}\frac{D_{ds}}{D_{s}}.$ (8)
The subscript SIE is used as a reminder that the Einstein radius is usually derived by fitting a singular isothermal ellipsoid to the lensing features. It is worth noticing however that the Einstein radius is very robust to the choice of lens model and only varies by 1-2% if a power law model is adopted instead (Bolton et al. 2008b).
Treu et al. (2006) introduced $f_{\rm SIE}$ as a proxy for the slope of the mass density profile, finding that it should be close to unity, as expected for an isothermal mass density profile. This can be understood qualitatively in terms of spherical profiles. Given a power law total mass density profile with three-dimensional form $\rho_{\rm tot}\propto r^{-\gamma}$, the mass enclosed within a two-dimensional circular aperture with radius $R$ is given by
$M(<R)=M_{\rm E}\left(\frac{R}{R_{\rm E}}\right)^{3-\gamma},$ (9)
where RE is the Einstein radius in kpc and ME is the mass within the Einstein radius. Following Schneider et al. (2006), it is easy to show that
$\sigma_{\rm SIE}^{2}R_{\rm E}=\frac{GM_{\rm E}}{\pi}.$ (10)
The virial theorem holds that
$\sigma_{1/2}^{2}R_{\rm eff}/2=\frac{GM(<R_{\rm eff}/2)}{c_{\rm e2}},$ (11)
where Reff is the effective radius and $c_{\rm e2}$ is a dimensionless structural parameter than depends on the three dimensional shape of the mass distribution, orbital structure, and spatial distribution of luminous tracers (Nipoti et al. 2008). Bolton et al. (2008b) finds $\langle\log c_{\rm e2}\rangle=0.53\pm 0.06$. In the following we will assume that ce2 is a constant, which is a good approximation. Residuals trends from the predicted relations would imply that ce2 is not exactly a constant, but that it depends, e.g., on galaxy mass, stellar velocity dispersion, and mass density profile slope.
By combining equations 9 to 11 one finds that
$f_{\rm SIE}^{2}=\frac{\pi}{c_{\rm e2}}\left(\frac{R_{\rm eff}/2}{R_{\rm E}}\right)^{2-\gamma}.$ (12)
In Figure 10 we compare the measured fSIE with the values predicted by Equation 12, using ce2 measured by Nipoti et al. (2008). Even though the data are noisy, we note that the SLACS-KWCI and TDCOSMO samples are very close to the predicted values. There is more scatter for the SL2S, which is in part expected owing to the fact that $R_{\rm eff}\sim R_{\rm E}/2$, and thus the factor to be elevated to the power of $2-\gamma_{\rm lens}$ is on average 4, introducing extrapolation errors. Furthermore, ce2 is not a constant and is expected to depend significantly on $\gamma$. Thus using the average value will be accurate for systems with $\gamma\sim 2$, similar to the SLACS sample and less so if $\gamma$ departs from 2.
Figure 10: Comparison between the measured $f_{\rm SIE}$ and that derived from Equation 12.
5.2 The lensing mass fundamental planes
Bolton et al. (2007) introduced the lensing mass fundamental plane (MFP; see also Bolton et al. 2008b; Auger et al. 2010) by analogy with the standard fundamental plane, i.e. replacing surface brightness with surface mass density:
$\log R_{\rm eff}=a\log\sigma_{1/2}+b\log\Sigma_{1/2}+c.$ (13)
This MFP could provide some useful insight. For a power law, the surface mass density within a projected radius is given by
$\Sigma=\Sigma_{c}\left(\frac{R}{R_{\rm E}}\right)^{1-\gamma},$ (14)
where $\Sigma_{c}$ is the lensing critical density
$\Sigma_{c}=\frac{M_{\rm E}}{\pi R_{\rm E}^{2}}=\frac{c^{2}D_{s}}{4\pi GD_{d}D_{ds}}.$ (15)
Thus, Equation 11 can be rewritten as
$\sigma_{1/2}^{2}=\frac{\pi G\Sigma_{1/2}R_{\rm eff}}{2c_{\rm e2}},$ (16)
where $\Sigma_{1/2}$ is the average surface mass density within half the effective radius. Taking the logarithm of both sides and sorting the variables yields
$\log R_{\rm eff}=2\log\sigma_{1/2}-\log\Sigma_{1/2}-\log\frac{\pi G}{2c_{\rm e2}}.$ (17)
The coefficients are found to be close to $a=2$ and $b=-1$ respectively, indicating that the lensing mass fundamental plane is much better described by this simple interpretation than the classic fundamental plane, which appears to be ”tilted” with respect to the virial prediction, requiring variations in M/L ratio with $\sigma$, generally attributed to changes in dark matter content, stellar population age, metallicity, and initial mass function (Treu et al. 2006, 2010; Auger et al. 2010). The intercept of course depends on the units. Consistent results, with coefficients in agreement with virial predictions, are obtained when the mass fundamental plane is measured using stellar dynamics fitted to integral-field stellar kinematics (e.g., Cappellari et al. 2006, 2013a; Zhu et al. 2024).
Previous studies (Bolton et al. 2007, 2008a; Auger et al. 2010) computed the lensing mass fundamental plane assuming isothermal mass profiles ($\gamma=2$). In Figure 11 we derive the MFP under the same assumption, using the SL2S sample, SLACS-KCWI sample, and the SLACS sample reported by Auger et al. (2010) based on SDSS velocity dispersions (SLACS-X). To fit the plane, we use Python package LtsFit (Cappellari et al. 2013b), a robust least squares regression with uncertainties in all dimensions and intrinsic scatter. We do not clip outliers but include all datapoints in the fit. Effective radii are fit in kpc, velocity dispersions in units of 250 $\mathrm{km\ s^{-1}}$, and surface mass densities in $\mathrm{10^{10}\ M_{\odot}\ kpc^{-2}}$. The samples follow the same MFP with coefficients consistent with those found by Auger et al. (2010). The slopes are close but not identical to those in Equation 17, indicating that the assumption of a constant virial coefficient is not exactly correct, but it is likely to have some dependence on mass or velocity dispersion, as elliptical galaxies are not completely scale invariant. We overplot the TDCOSMO sample for illustration, showing them to lie along the plane across the sample space occupied by the SL2S and SLACS samples. The model indicates no intrinsic scatter to the combined sample.
Figure 11: Isothermal Mass Plane (MP) $\mathrm{\log R_{eff}=a\log\sigma_{1/2}+b\log\Sigma_{1/2,iso}+c}$. Plane is fitted to the combined SL2S, SLACS/KCWI, and SLACS-X samples. TDCOSMO sample is overplotted. Here $\mathrm{\Sigma_{1/2,iso}=M_{R_{eff}/2,iso}/\pi(R_{eff}/2)^{2}}$ is the average lensing surface mass density within half the effective radius assuming the mass density profile is isothermal (i.e., $\gamma_{lens}=2$). $\sigma_{1/2}$ is given in units of 250 $\mathrm{km\ s^{-1}}$, $\mathrm{M_{1/2}}$ in units of $10^{10}\ \mathrm{M_{\odot}}$, and $\mathrm{R_{eff}}$ in kpc. As a test, we fitted all combinations of subsamples, and the coefficients are the same within the errors. The dotted line is the one-to-one line.
We are also in a position to study for the first time the MFP for a generic power law profile, using the measurements of $\gamma=\gamma_{\rm lens}$ that are now available. The uncertainty on $\Sigma_{1/2}$ is given by
$\delta\log\Sigma_{1/2}=\sqrt{(1-\gamma_{\rm lens})^{2}\left(\frac{\delta R_{\rm eff}}{R_{\rm eff}\ln 10}\right)^{2}+\delta\gamma^{2}\left(\log\frac{R_{\rm eff}}{2R_{E}}\right)^{2}}$ (18)
The resulting MFP is shown in Figure 12. As expected, the errors are larger than in the isothermal version, especially for the SL2S sample. For this sample, $\delta\gamma_{\rm lens}$ is generally larger than for TDCOSMO and SLACS due to the lower SNR of the imaging data. Furthermore, the Einstein radius is typically several times the effective radius, requiring more extrapolation than for the other samples. This plane is consistent with no scatter within its error bars as well, but in Figure 12 we report the best fit, which includes 1% intrinsic scatter.
Both planes are well defined, and there are no significant differences between the sub-samples. We quantify any residual differences between the sub-samples in the following way. Since the sub-samples, except SLACS-X, are too small to robustly determine the slopes, we keep the slopes $a$ and $b$ fixed to the values obtained for the combined sample and then we fit for the intercept $c$ of each subsample. In the isothermal case, the intercept is $0.451\pm 0.019$, $0.463\pm 0.027$, $0.456\pm 0.027$, and $0.445\pm 0.015$, respectively for the combined, SL2S, SLACS-KCWI, and SLACS-X samples. They are consistent within the errors. The error on the intercept is largest for the SL2S sample and corresponds to $\pm$5.8in Reff and ±$\pm 4.2\%$ in $\sigma_{1/2}$. For the non-isothermal case, the intercept is $0.560\pm 0.020$, $0.579\pm 0.029$, $0.561\pm 0.021$, and $0.548\pm 0.022$ for the samples in the same order as above. Again, they are consistent within the errors. The error is largest for the SL2S sample and corresponds to $\pm$6.9in Reff and ±$\pm 5.3\%$ in $\sigma_{1/2}$.
Figure 12: Mass plane (MP) $\mathrm{\log R_{eff}=a\log\sigma_{1/2}+b\log\Sigma_{1/2}+c}$. The plane is fitted to the combined SL2S, SLACS/KCWI, and SLACS-X samples. TDCOSMO sample is overplotted. Here $\mathrm{\Sigma_{1/2}=M_{R_{eff}/2}/\pi(R_{eff}/2)^{2}}$ is the average lensing surface mass density within half the effective radius. We have updated the SLACS-X lens models to those from Tan et al. (2024) with free power law slopes to compare with SL2S and SLACS/KCWI lens models. $\sigma_{1/2}$ is given in units of 250 $\mathrm{km\ s^{-1}}$, $\mathrm{M_{1/2}}$ in units of $10^{10}\ \mathrm{M_{\odot}}$, and $\mathrm{R_{eff}}$ in kpc. The dotted line is the one-to-one line.
In conclusion, the MFP analysis shows no evidence of structural differences between the SL2S, SLACS, and TDCOSMO samples. This does not mean that there is no evolution in the characteristics of the galaxies. For example, in Figures 11 and 12 the SL2S galaxies tend to have smaller effective radii, generally consistent with the finding that early type galaxies have smaller sizes at high redshift that in the local universe for a given stellar mass (e.g., van Dokkum et al. 2010). Sonnenfeld et al. (2015) found that the fraction of dark matter within the inner 5 kpc increases with redshift, and Sheu et al. (2025, submitted) also found mild evidence for evolution. However, this work shows that galaxies evolve within the lensing mass plane. An example of a mechanism that moves galaxies along the lensing mass plane is the ’envelope accretion’ recently considered by Nipoti (2025) in addition to the standard major and minor mergers. Envelope accretion describes the case in which diffuse satellites are completely disrupted in the galaxy outskirts. As show in Figure 13, tantalizingly, the model proposed by Nipoti (2025) predicts that envelope accretion makes galaxies evolve along the lensing mass plane, almost exactly. Similar results are obtained for the generalized mass plane with free mass density slopes. A full comparison with galaxy evolution models is beyond the scope of this paper and is left for future work.
Figure 13: Evolution of a reference galaxy in the lens mass fundamental plane according to the ”envelope accretion” model by Nipoti (2025). The initial (red symbol) and final (blue symbol) positions of the galaxy in the plane are calculated using the model described in section 6.2 of Nipoti (2025), in which the galaxy grows by a factor of two in stellar mass.
6 Summary
We have applied improved methodologies and clean stellar template libraries (Knabel et al. 2025) to obtain new stellar velocity dispersions using spectra obtain at the Keck and VLT Telescopes for the lens galaxies in the SL2S sample (Gavazzi et al. 2012; Ruff et al. 2011; Sonnenfeld et al. 2013a, b, 2015, Sheu et al. 2025, submitted). The SL2S lens sample is very well matched in terms of deflector redshift and Einstein radii to quasar lenses used by the TDCOSMO collaboration for time delay cosmography (Millon et al. 2020), and it can be used as an external dataset (Birrer et al. 2020) to improve constraints on the Hubble constant and other cosmological parameters (Birrer & Treu 2021), provided stellar velocity dispersions are accurately measured. We use these new measurements to investigate the lens mass fundamental plane (Bolton et al. 2007, 2008a; Auger et al. 2010) as a tool to compare the internal structure of massive galaxies. Our main results can be summarized as follows:
  • The new measurements are on average consistent with the old measurements, but there is significant scatter on an individual system basis.
  • Systematic errors are dominated by the choice of stellar templates and are of the order of 2-3%, depending on the SNR of the spectra.
  • Stellar templates introduce positive covariance between velocity dispersion measurements across the galaxy samples. The covariance is at the level of 1-2%, depending on spectral quality, for galaxies observed with the same instrumental setup. It is significantly smaller (sub-percent) for spectra obtained with different instrumental setups.
  • We derive a simple formula relating the slope of the mass density profile to the observed stellar velocity dispersion, Einstein radius, and effective radius under the assumption of a constant virial coefficient, and show that it is consistent within the errors with the observations.
  • The SL2S, SLACS, and TDCOSMO galaxies follow a tight lens mass fundamental plane, consistent within the errors with the one derived for the SLACS galaxies alone.
  • We introduce a new version of the lens mass fundamental plane, relaxing the assumption of isothermal mass density profiles. We find that this new version is also well defined, albeit with larger error bars as expected, and the three samples are mutually consistent within $1\sigma$ uncertainty.
In terms of big picture, two major conclusions of this study are the following. First, with the new kinematic measurements and systematic error estimates, the SL2S galaxies with the best spectra can be used as external datasets for time delay cosmography. Second, the internal mass structure of massive elliptical galaxies evolves with redshift. However, this evolution happens within the lensing mass plane, which sets tight constraints on their evolutionary mechanism, such as dry mergers (Nipoti et al. 2009) or envelope accretion (Nipoti 2025).
Acknowledgements.
The first two authors should be regarded as joint first authors. We are grateful to William Sheu and our friends and colleagues of the TDCOSMO collaboration for many useful discussions on these topics. TT acknowledges support by NSF through grants NSF-AST-1906976 and NSF-AST-1836016, and from the Moore Foundation through grant 8548. AJS received support from NASA through the STScI grants HST-GO-16773 and JWST-GO-2974.
Table 1. For each lens we list the instruments used, SNR per Å of the spectra, measured instrumental resolution, BIC weighted and equally weighted velocity dispersions, and associated statistical and systematic uncertainties. The BIC determined weights used for each galaxy are reported in Table 2. The systematic uncertainties are Bessel corrected. The last column shows the relative error associated with uncertainties in the measurement of instrumental resolution. For galaxies with two spectra of sufficient spectral resolution, we report both values. The one with the highest SNR is listed first, and it is our fiducial measurement.
Name Instrument S/N $\sigma_{\text{inst}}$ $\overline{\sigma}_{\text{BIC\_w}}$ $\delta\overline{\sigma}_{\text{BIC\_w}}$ $\overline{\sigma}_{\text{eq\_w}}$ $\delta\overline{\sigma}_{\text{eq\_w}}$ $\Delta_{B}\overline{\sigma}_{\text{BIC\_w, eq\_w}}$ $\frac{\delta\sigma_{\text{inst}}}{\sigma_{\text{inst}}}(\frac{\sigma_{\text{inst}}}{\sigma})^{2}$
-1) (km s-1) (km s-1) (km s-1) (km s-1) (km s-1) (km s-1)
SL2SJ020833-071414 LRIS 14 64 $\pm$ 6 252 19 252 19 3 0.006
SL2SJ021206-075528 LRIS 23 52 $\pm$ 7 236 7 235 8 2 0.006
SL2SJ021247-055552 XSHOOTER 15 32 $\pm$ 2 256 10 254 10 4 0.001
SL2SJ021325-074355 LRIS 4 97 $\pm$ 8 344 24 352 24 13 0.007
SL2SJ021411-040502 XSHOOTER 23 32 $\pm$ 2 209 7 213 7 7 0.001
LRIS 10 59 $\pm$ 6 250 15 254 15 7 0.006
SL2SJ021737-051329 LRIS 9 97 $\pm$ 8 253 20 258 19 8 0.012
LRIS 8 68 $\pm$ 7 248 14 249 14 2 0.008
SL2SJ021801-080247 XSHOOTER 5 31 $\pm$ 2 254 28 255 28 3 0.001
SL2SJ021902-082934 LRIS 17 99 $\pm$ 4 285 7 284 7 7 0.005
SL2SJ022046-094927 XSHOOTER 8 32 $\pm$ 2 187 16 188 16 1 0.002
SL2SJ022056-063934 LRIS 18 97 $\pm$ 8 246 11 245 11 8 0.013
SL2SJ022346-053418 LRIS 15 64 $\pm$ 6 298 16 299 16 7 0.004
SL2SJ022357-065142 LRIS 20 59 $\pm$ 4 254 16 253 16 4 0.003
LRIS 18 68 $\pm$ 7 258 20 270 21 19 0.007
SL2SJ022610-042011 LRIS 20 94 $\pm$ 6 310 13 311 13 6 0.006
LRIS 13 93 $\pm$ 9 287 11 284 11 5 0.01
SL2SJ022648-040610 XSHOOTER 31 32 $\pm$ 2 337 9 343 8 8 0.001
SL2SJ022648-090421 LRIS 16 97 $\pm$ 8 326 8 323 8 12 0.007
LRIS 7 68 $\pm$ 7 350 24 353 23 5 0.004
SL2SJ023251-040823 XSHOOTER 35 30 $\pm$ 1 208 5 208 5 1 0.001
LRIS 16 93 $\pm$ 9 240 9 237 9 4 0.015
SL2SJ084847-035103 XSHOOTER 13 33 $\pm$ 2 198 8 198 8 2 0.002
SL2SJ084909-041226 DEIMOS 17 61 $\pm$ 6 338 14 340 14 5 0.003
SL2SJ084934-043352 LRIS 19 56 $\pm$ 2 221 11 221 11 0 0.003
SL2SJ084959-025142 XSHOOTER 65 33 $\pm$ 0 208 4 208 4 1 0.0
SL2SJ085019-034710 LRIS 20 58 $\pm$ 1 234 12 232 12 10 0.001
SL2SJ085540-014730 LRIS 18 56 $\pm$ 2 196 11 195 10 1 0.003
SL2SJ085559-040917 LRIS 27 51 $\pm$ 7 275 8 279 8 11 0.005
SL2SJ085826-014300 LRIS 8 64 $\pm$ 6 240 14 241 14 2 0.007
SL2SJ090106-025906 XSHOOTER 5 33 $\pm$ 2 224 25 219 25 27 0.001
SL2SJ090407-005952 XSHOOTER 19 33 $\pm$ 1 232 9 233 9 0 0.001
SL2SJ095921+020638 XSHOOTER 17 33 $\pm$ 1 200 8 199 8 0 0.001
SL2SJ135847+545913 LRIS 22 62 $\pm$ 9 256 9 256 9 1 0.008
SL2SJ135949+553550 LRIS 9 65 $\pm$ 6 221 14 222 15 1 0.008
LRIS 7 68 $\pm$ 7 219 20 219 19 0 0.01
SL2SJ140123+555705 LRIS 9 85 $\pm$ 13 379 18 371 18 13 0.008
SL2SJ140156+554446 LRIS 28 64 $\pm$ 5 276 9 275 9 3 0.004
SL2SJ140454+520024 LRIS 31 63 $\pm$ 6 336 9 342 9 15 0.003
SL2SJ140546+524311 LRIS 19 61 $\pm$ 8 258 8 258 8 2 0.007
SL2SJ140614+520253 LRIS 8 85 $\pm$ 13 268 14 272 14 6 0.015
SL2SJ140650+522619 LRIS 12 65 $\pm$ 6 270 11 267 11 6 0.005
SL2SJ142031+525822 LRIS 19 70 $\pm$ 3 218 11 216 11 7 0.005
SL2SJ142059+563007 LRIS 20 63 $\pm$ 8 244 16 247 16 7 0.008
LRIS 15 68 $\pm$ 7 250 10 250 10 0 0.008
SL2SJ220329+020518 LRIS 29 66 $\pm$ 2 203 10 208 10 10 0.002
SL2SJ220506+014703 XSHOOTER 31 32 $\pm$ 2 319 7 309 7 14 0.001
SL2SJ220629+005728 LRIS 4 97 $\pm$ 8 281 27 286 27 9 0.01
SL2SJ221326-000946 XSHOOTER 30 31 $\pm$ 2 194 12 193 11 1 0.002
LRIS 23 72 $\pm$ 7 162 14 165 13 5 0.019
SL2SJ221407-180712 LRIS 5 97 $\pm$ 8 200 16 208 16 13 0.019
SL2SJ221852+014038 LRIS 19 64 $\pm$ 3 292 11 291 11 4 0.002
SL2SJ221929-001743 LRIS 26 55 $\pm$ 3 242 6 241 7 0 0.002
LRIS 32 93 $\pm$ 9 239 11 242 11 6 0.015
SL2SJ222012+010606 DEIMOS 17 79 $\pm$ 1 248 15 248 14 1 0.002
SL2SJ222148+011542 DEIMOS 29 74 $\pm$ 5 277 10 283 10 8 0.005
SL2SJ222217+001202 LRIS 11 63 $\pm$ 4 198 20 196 20 3 0.006
LRIS 8 68 $\pm$ 7 190 28 195 29 8 0.013
Table 2: We report the BIC-weighted average statistical uncertainty, the average systematic uncertainty (both Bessel corrected and uncorrected), the average amplitude of the off-diagonal terms of the covariance matrix between elements of the sample (both Bessel corrected and uncorrected) for the whole SL2S sample and subsamples grouped by instrument and SNR. The BIC determined weights given to the two libraries for each subsample, and the number of galaxies in each subsample is also provided.
Sample $<\delta\bar{\sigma}/\bar{\sigma}>$ $<\Delta\bar{\sigma}/\bar{\sigma}>$ $\sqrt{<C_{i,j}\bar{\sigma}/\bar{\sigma_{i}}\bar{\sigma_{j}}>_{i\neq j}}$ $<\Delta_{B}\bar{\sigma}/\bar{\sigma}>$ $\sqrt{<C_{B,i,j}\bar{\sigma}/\bar{\sigma_{i}}\bar{\sigma_{j}}>_{i\neq j}}$ Indo-US MILES Num. of Obj.
SL2S S/N ¿ 20 3.47% 0.66% 0.22% 2.14% 0.72% 0.05 0.95 15
SL2S S/N (13, 20] 4.60% 1.07% 0.65% 1.70% 1.04% 0.27 0.73 18
SL2S S/N $\leq$ 13 7.40% 1.68% 0.90% 3.07% 1.65% 0.81 0.19 14
SL2S All 5.05% 1.38% 0.67% 2.22% 1.08% 0.26 0.74 47
SL2S-LRIS S/N ¿ 20 3.73% 1.39% 1.07% 2.38% 1.83% 0.78 0.22 8
SL2S-LRIS S/N (13, 20] 4.68% 1.42% 0.97% 2.09% 1.42% 0.37 0.63 13
SL2S-LRIS S/N $\leq$ 13 6.90% 0.91% 0.43% 2.73% 1.30% 0.94 0.06 10
SL2S-LRIS All 5.13% 1.63% 1.02% 2.36% 1.48% 0.62 0.38 31
SL2S-Other S/N ¿ 20 3.16% 0.003% 0.001% 1.97% 0.45% 0.001 0.999 7
SL2S-Other S/N (13, 20] 4.38% 0.40% 0.04% 0.71% 0.06% 0.19 0.81 5
SL2S-Other S/N $\leq$ 13 8.68 % 2.49% 1.55% 3.69% 2.30% 0.35 0.65 4
SL2S-Other All 4.94% 0.10% 0.02% 1.96% 0.35% 0.001 0.999 16
Table 3: Values used in scaling relation calculations and figures. For each dataset (S2LS, SLACS/KCWI, TDCOSMO), we report object name, deflector and source redshift, effective radius, Einstein radius, stellar velocity dispersion corrected to an equivalent aperture of radius equal to half the effective radius, and slope of the mass density profile as inferred from lensing. References are: SL2S (Sonnenfeld et al. 2013a, Sheu et al. 2025; submitted); SLACS(Auger et al. 2010; Tan et al. 2024; Knabel et al. 2024); TDCOSMO (Birrer et al. 2020, and TDCOSMO collaboration 2025, in prep.)
Name $\mathrm{z_{d}}$ $\mathrm{z_{s}}$ $\mathrm{R_{eff}}$ $\mathrm{\theta_{E}}$ $\mathrm{\sigma_{1/2}}$ $\mathrm{\gamma_{lens}}$
[…]
SL2SJ021411-040502 0.609 1.88 0.42 1.34 $\pm$ 0.03 219 $\pm$ 10 2.39 $\pm$ 0.2
SL2SJ021902-082934 0.389 2.15 0.39 1.30 $\pm$ 0.03 303 $\pm$ 10 2.70 $\pm$ 0.11
SL2SJ022610-042011 0.494 1.23 0.39 1.15 $\pm$ 0.02 324 $\pm$ 15 2.05 $\pm$ 0.15
SL2SJ023251-040823 0.352 2.34 0.32 1.02 $\pm$ 0.02 221 $\pm$ 5 2.38 $\pm$ 0.18
SL2SJ084909-041226 0.722 1.54 0.24 1.14 $\pm$ 0.03 363 $\pm$ 16 2.07 $\pm$ 0.15
SL2SJ084959-025142 0.274 2.09 1.54 1.19 $\pm$ 0.01 207 $\pm$ 3 2.24 $\pm$ 0.30
SL2SJ085540-014730 0.365 3.39 0.47 0.96 $\pm$ 0.02 203 $\pm$ 11 2.25 $\pm$ 0.11
SL2SJ090407-005952 0.611 2.36 0.71 1.41 $\pm$ 0.03 238 $\pm$ 9 1.98 $\pm$ 0.15
SL2SJ095921+020638 0.552 3.35 0.41 0.71 $\pm$ 0.02 209 $\pm$ 8 2.12 $\pm$ 0.18
SL2SJ140454+520024 0.456 1.59 0.78 2.39 $\pm$ 0.05 360 $\pm$ 18 2.49 $\pm$ 0.21
SL2SJ140546+524311 0.526 3.01 0.48 1.46 $\pm$ 0.03 270 $\pm$ 9 2.23 $\pm$ 0.17
SL2SJ142059+563007 0.483 3.12 1.79 1.39 $\pm$ 0.01 283 $\pm$ 13 2.06 $\pm$ 0.22
SL2SJ220329+020518 0.400 2.15 0.99 1.68 $\pm$ 0.03 217 $\pm$ 15 1.94 $\pm$ 0.08
SL2SJ220506+014703 0.476 2.53 0.40 1.81 $\pm$ 0.04 333 $\pm$ 16 1.92 $\pm$ 0.14
SL2SJ222148+011542 0.325 2.35 0.53 1.27 $\pm$ 0.03 290 $\pm$ 13 2.14 $\pm$ 0.13
SDSSJ002907.77-005550.5 0.227 0.931 2.30 0.95 $\pm$ 0.02 207 $\pm$ 7 2.47 $\pm$ 0.26
SDSSJ003753.21-094220.1 0.195 0.632 2.30 1.48 $\pm$ 0.03 279 $\pm$ 7 2.38 $\pm$ 0.11
SDSSJ033012.14-002051.9 0.351 1.071 1.26 1.09 $\pm$ 0.02 258 $\pm$ 12 2.14 $\pm$ 0.17
SDSSJ111250.60+082610.4 0.273 0.629 1.55 1.50 $\pm$ 0.03 274 $\pm$ 7 1.92 $\pm$ 0.18
SDSSJ120444.07+035806.4 0.164 0.631 1.63 1.30 $\pm$ 0.03 262 $\pm$ 5 2.02 $\pm$ 0.06
SDSSJ125028.26+052349.1 0.232 0.795 1.86 1.12 $\pm$ 0.02 243 $\pm$ 6 2.18 $\pm$ 0.09
SDSSJ130613.65+060022.1 0.173 0.472 2.08 1.31 $\pm$ 0.03 230 $\pm$ 7 1.97 $\pm$ 0.07
SDSSJ140228.21+632133.5 0.205 0.481 2.65 1.35 $\pm$ 0.03 285 $\pm$ 7 2.24 $\pm$ 0.15
SDSSJ153150.07-010545.7 0.160 0.744 2.73 1.71 $\pm$ 0.03 273 $\pm$ 9 2.07 $\pm$ 0.25
SDSSJ153812.92+581709.8 0.143 0.531 1.45 1.00 $\pm$ 0.02 236 $\pm$ 5 2.27 $\pm$ 0.12
SDSSJ162132.99+393144.6 0.245 0.602 2.30 1.27 $\pm$ 0.03 255 $\pm$ 10 2.02 $\pm$ 0.07
SDSSJ162746.45-005357.6 0.208 0.524 2.02 1.22 $\pm$ 0.02 262 $\pm$ 10 1.93 $\pm$ 0.14
SDSSJ163028.16+452036.3 0.248 0.793 2.01 1.79 $\pm$ 0.04 280 $\pm$ 7 1.92 $\pm$ 0.09
SDSSJ230321.72+142217.9 0.155 0.517 3.46 1.62 $\pm$ 0.03 259 $\pm$ 9 1.94 $\pm$ 0.08
B1608+656 0.630 1.394 0.59 0.81$\pm$0.05 266$\pm$16 2.08$\pm$0.11
DES0408-5354 0.597 2.375 1.94 1.92$\pm$0.10 222$\pm$26 1.90$\pm$0.10
HE0435-1223 0.455 1.693 1.80 1.22$\pm$0.08 214$\pm$15 1.93$\pm$0.10
PG1115+080 0.311 1.722 0.45 1.08$\pm$0.06 292$\pm$26 2.17$\pm$0.12
RXJ1131-1231 0.295 0.654 1.91 1.63$\pm$0.08 313$\pm$20 1.95$\pm$0.11
SDSS1206+4332 0.745 1.789 0.29 1.25$\pm$0.06 310$\pm$32 1.95$\pm$0.11
WFI2033-4723 0.658 1.662 1.97 0.94$\pm$0.05 250$\pm$19 1.95$\pm$0.10
WGD2038-4008 0.228 0.777 2.22 1.38$\pm$0.07 286$\pm$19 2.3$\pm$0.12